# Curious Causality-Seeking Agents Learn Meta Causal World

### v2

@@ Abstract
When building a world model, a common assumption is that the environment has a single, unchanging underlying causal rule, like applying Newton’s laws to every situation. In reality, what appears as a drifting causal mechanism is often the manifestation of a fixed underlying mechanism seen through a narrow observational window. This brings about a problem that, when building a world model, even subtle shifts in policy or environment states can alter the very observed causal mechanisms. In this work, we introduce the Meta-Causal Graph as world models, a minimal unified representation that efficiently encodes the transformation rules governing how causal structures shift across different latent world states. A single Meta-Causal Graph is composed of multiple causal subgraphs, each triggered by meta state, which is in the latent state space. Building on this representation, we introduce a Causality-Seeking Agent whose objectives are to (1) identify the meta states that trigger each subgraph, (2) discover the corresponding causal relationships by agent curiosity-driven intervention policy, and (3) iteratively refine the Meta-Causal Graph through ongoing curiosity-driven exploration and agent experiences. Experiments on both synthetic tasks and a challenging robot arm manipulation task demonstrate that our method robustly captures shifts in causal dynamics and generalizes effectively to previously unseen contexts.
@@ 1 Introduction
To identify the underlying world rules, traditional causal models rely predominantly on observational data (pmlr-v216-lippe23a), implicitly assuming a fixed causal structure. In open-ended environments, although the ground-truth causal laws remain unchanged, the observational data is collected through local, context-limited exploration. Under conditions of partial observability, this invariance becomes obscured within a limited observational window, giving rise to the illusion of shifting causal relationships—a rule valid in one context may break down in another. For instance, the "push → open" relation holds for an unlocked door, but this relation vanishes when the door is locked. In this case, the same action produces no effect, leaving the true causal mechanism ambiguous. Traditional causal modeling in such scenarios faces two main challenges: (1) methods that learn a single uniform (time series-based) causal graph neglect these context-dependent changes (NEURIPS2023_65496a49), and (2) lack of an active curiosity-driven exploration strategy, which makes it struggle to adapt to discover the global causal rules in rich, unseen variations of potential open-ended worlds (huang2020causal; hwang2024fine).
@@ 2 Preliminaries
Causal graph
A causal subgraph is formalized as a directed acyclic graph (DAG) $\mathcal{G}=([p],E)$, where the vertex set $[p]=\{1,2,\dots,p\}$ indexes a collection of random variables $X=\{X_{i}\}_{i=1}^{p}$ with joint density $f(X)$. The edge set $E\subseteq[p]\times[p]$ encodes direct causal relationships between variables. For each node $i\in[p]$, we write $\mathrm{Pa}_{\mathcal{G}}(i)$ for its set of parents. An edge $i\to j\in E$ indicates that $X_{i}$ exerts a direct causal influence on $X_{j}$.
Intervention
@@
During world-model learning, we perform a sequence of environment interventions to reveal causal structure. To better formalize the intervention influence, in this paper, we introduce the notion of an intervention target. At intervention step $k$, we select a single variable $X_{i}$ as the target, which we denote by the intervention target $I_{k}=\{i\}$, where $i\in[p]$. Applying an intervention $\mathrm{do}(X_{I_{k}}=x^{\prime})$ removes all incoming causal links into $X_{i}$. Concretely, if $\mathcal{G}$ is the original DAG, the intervention graph $\mathcal{G}^{(k)}$ is obtained by removing every edge of the form $j\to i,\quad\forall\,j\in\mathrm{Pa}_{\mathcal{G}}(i),$ from $\mathcal{G}$. This modified graph accurately reflects the altered generative process under the do-operation on $X_{i}$.
@@
[…]
@@ 3 Meta-Causal Graph: Definition and Identifiability
In this section, we present our Meta-Causal Graph for a world formulation. We begin by defining the Meta-Causal Graph, which is a compact causal graph comprising multiple context-specific causal subgraphs, varying from latent meta states in Section 3.1. Then, we show the identifiability of meta states in Section 3.2 and their indicated causal subgraphs in Section 3.3 by interventions.
@@ 3.1 Meta-Causal Graph
The Meta-Causal Graph $\mathcal{MG}$ consists of a collection of causal subgraphs $\mathcal{MG}\;=\;\{\mathcal{G}_{u}\}_{u\in U},$ where each $\mathcal{G}_{u}$ corresponds to a distinct meta state $u\in U$. The causality skeleton matrix $M_{u}$ is a binary matrix where $M_{u}[i,j]=1$ indicates that variable $i$ is a parent of variable $j$ (i.e., $i\in Pa(j)$) in the causal graph $\mathcal{G}_{u}$. A state-to-meta state mapping $C:\mathcal{X}\;\to\;U$ assigns each system observation $\mathbf{x}\in\mathcal{X}$ to its active meta state $u=C(\mathbf{x})$. We call $\mathcal{MG}=(\{\mathcal{G}_{u}\},C)$ a Meta-Causal Graph if the following conditions are satisfied:
@@
[…]
@@ 3.3.1 Causal Subgraph Identification
We employ multiple interventions to determine the causal structure of environment, represented as intervention target indicator set =$\mathcal{I}=\{I_{k}\}_{k=1}^{K}$, where $I_{k}\subset\{[p]\}$ comprises the indices of intervened variables selected from state variable set $X=\{X_{i}\}_{i\in[p]}$ at intervention step $k$. These intervention sets are collected from historical data or agent interactions. We establish sufficient conditions ensuring causal graph edge identifiability through appropriate interventional selection.
@@
4 Curious Causality-Seeking Agents for Meta Causal World Modeling
@@
We propose a novel framework to learn the Meta-Causal Graph by a Causality-Seeking Agent. Our approach consists of three main components: (1) Meta-Causal Graph discovery from agent experience, (2) curiosity-driven interventional exploration for seeking Meta-Causal Graph in an unknown world and (3) complete world model learning and updating. In this section, we detail each component and describe how they work together to enable discovery of causal structures.
4.1 Meta-Causal Graph Learning from Experience
We learn causal subgraphs under latent meta states from agent experience. Following FCDL’s pioneering use of the VQ-VAE architecture for causal-graph discovery (hwang2024fine), we adopt the original VQ-VAE framework of van den Oord et al. (van2017neural) to learn the causal subgraph for each meta-state. Specifically, the process contains two steps: (1) identifying the latent meta state, which activates a specific causal subgraph, and (2) learning the subgraph from the agent’s explored experience data.
Identifying Latent Meta State: We embed the observed state and assign it to a corresponding meta state $u\in U$ by vector quantization as follows. We define the meta state assignment as $C(x)\;=\;\arg\min_{u\in U}\bigl{\lVert}E(x)-e_{u}\bigr{\rVert}_{2}^{2},$ where $E\colon\mathcal{X}\to\mathbb{R}^{d}$ is an encoder mapping the observed state $x$ to a $d$-dimensional embedding, and $\{e_{u}\}_{u\in U}\subset\mathbb{R}^{d}$ is a learnable codebook of prototype embeddings, each representing a distinct meta state.
Subgraph Learning: After obtaining the embedding $e_{u}$ for a given state, we employ a decoder network $D:\mathbb{R}^{d}\;\to\;[0,1]^{p\times p}$ to predict a probability matrix $\hat{M}_{u}$, which estimates the underlying causal skeleton matrix $M_{u}$ of $\mathcal{G}_{u}$. Each entry $\hat{M}_{u}[i,j]$ denotes the probability that the directed edge $i\to j$ is present in meta state $u$. We then sample each $M_{u}[i,j]$ from $\mathrm{Bernoulli}(\hat{M}_{u}[i,j])$ using the Gumbel-Softmax reparameterization (jang2016categorical; maddison2016concrete), enabling end-to-end gradient-based learning of the discrete structure. To update the codebook, we use the following update rule:
$\mathcal{L}_{\text{quantization}}=\|\text{sg}(E_{\phi}(X))-e_{u}\|_{2}^{2}+\beta\|\text{sg}(e_{u})-e_{u}\|_{2}^{2},$
[…]
4.2 Curious Reward of Interventional Exploration for Seeking Causality in Unknown World
Accurate identification of the Meta-Causal Graph cannot rely solely on passive agent experience. First, causal structures inferred from observational trajectories alone may misorient edges. Second, an agent’s experiences cover only a subset of possible world states, leaving portions of the causal graph unobserved. To efficiently explore the state space and discover transitions between different meta states, we design a curiosity-driven reward function:
$R(\text{do}(X_{i}^{t}=x^{\prime}_{i}))=\sum_{i,j\in[p]}H(\hat{M}_{C(X^{t})[i,j]}),$ (2)
where $H$ is the Shannon entropy of the posterior distribution over potential causal relationships given the current state $X^{t}$. This reward function encourages the agent to explore states in which the underlying causal structure remains uncertain.
To validate and refine the learned causal structures, we employ interventional verification to directly test whether specific variables causally influence others under different state conditions. For each variable $X_{i}$ in our system, we perform interventions denoted as $\text{do}(X_{i}=x^{\prime}_{i})$, where we intervene $X_{i}$ by setting it to a specific value $x^{\prime}_{i}$ and observe the consequent changes in other variables. The causal effect of $X_{i}$ on $X_{j}$ is then estimated as:
$\Delta_{ij}=\log P(X_{j}^{t+1}|X^{t},\text{do}(X_{i}=x^{\prime}_{i}))-\log P(X_{j}^{t+1}|X^{t}),$
where $\Delta_{ij}$ is the difference in log probability between the intervened and the observational distribution of $X_{j}$. We incorporate the causal effect estimation into the objective function as follows:
$\mathcal{L}_{\text{mask}}=-\lambda_{1}\sum_{\{(i,j)|\Delta_{ij}>0\}}\log\hat{M}_{ij}+\lambda_{2}\sum_{\{(i,j)|\Delta_{ij}<0\}}\log(\hat{M}_{ij}),$
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters.
@@ 4.3 Transition Probability Learning
MCG (ours) 63.18$\pm$13.94 50.47$\pm$9.87 50.04$\pm$8.56 51.99$\pm$6.58 49.78$\pm$4.11 49.69$\pm$5.14
MCG (wo mask loss) 58.18$\pm$14.51 51.70$\pm$8.58 46.04$\pm$8.56 47.75$\pm$7.52 46.19$\pm$5.87 47.94$\pm$6.60
MCG (wo intervention verification) 48.28$\pm$8.15 43.48$\pm$5.05 46.54$\pm$3.72 50.36$\pm$8.05 50.55$\pm$6.83 48.72$\pm$4.42
GNN 36.29$\pm$3.45 25.80$\pm$3.48 21.58$\pm$3.44 29.22$\pm$3.39 23.28$\pm$4.98 20.53$\pm$6.96
MLP 31.11$\pm$1.69 30.44$\pm$2.28 32.39$\pm$1.76 28.66$\pm$3.65 26.52$\pm$4.26 24.15$\pm$4.17
NCD 41.60$\pm$5.08 37.47$\pm$2.13 42.27$\pm$1.82 40.04$\pm$6.21 37.47$\pm$2.98 41.19$\pm$1.66
FCDL 57.82$\pm$9.90 49.29$\pm$8.90 47.70$\pm$6.68 50.66$\pm$10.10 48.81$\pm$8.91 48.05$\pm$5.86
Modular 26.53$\pm$3.45 24.73$\pm$5.61 26.73$\pm$8.31 25.24$\pm$4.68 24.94$\pm$4.81 25.09$\pm$5.91
NPS 40.56$\pm$4.61 26.81$\pm$4.37 23.02$\pm$4.27 38.73$\pm$2.63 27.69$\pm$4.28 24.45$\pm$3.84
CDL 35.59$\pm$1.85 35.82$\pm$1.40 42.22$\pm$1.39 34.90$\pm$1.59 36.52$\pm$1.72 42.06$\pm$1.29
GRADER 37.93$\pm$1.06 38.94$\pm$1.63 45.74$\pm$2.25 36.82$\pm$3.12 37.41$\pm$2.84 43.48$\pm$4.14
Oracle 33.87$\pm$1.34 36.48$\pm$1.80 42.47$\pm$0.75 34.63$\pm$1.78 38.31$\pm$2.48 42.87$\pm$2.08
@@
1: Initialize encoder parameters $\phi$, decoder parameters $\theta$, and embeddings $\{e_{u}\}_{u\in U}$
@@
3:  Compute embedding assignment: $u\leftarrow\operatorname*{arg\,min}_{u\in U}\|E(X)-e_{u}\|_{2}^{2}$
@@
6:  Select intervention $\text{do}(X_{i}=x^{\prime}_{i})$ maximizing reward $R(\text{do}(X_{i}=x^{\prime}_{i}))$ via Eq. (2)

### v3

# Curious Causality-Seeking Agents in Open-ended Worlds

@@ Abstract
When building a world model, a common assumption is that the environment has a single, unchanging underlying causal rule, like applying Newton’s laws to every situation. However, in truly open-ended environments, the apparent causal mechanism may drift over time because the agent continually encounters novel contexts and operates within a limited observational window. This brings about a problem that, when building a world model, even subtle shifts in policy or environment states can alter the very observed causal mechanisms. In this work, we introduce the Meta-Causal Graph as world models for open-ended environments, a minimal unified representation that efficiently encodes the transformation rules governing how causal structures shift across different latent world states. A single Meta-Causal Graph is composed of multiple causal subgraphs, each triggered by meta state, which is in the latent state space. Building on this representation, we introduce a Causality-Seeking Agent whose objectives are to (1) identify the meta states that trigger each subgraph, (2) discover the corresponding causal relationships by agent curiosity-driven intervention policy, and (3) iteratively refine the Meta-Causal Graph through ongoing curiosity-driven exploration and agent experiences. Experiments on both synthetic tasks and a challenging robot arm manipulation task demonstrate that our method robustly captures shifts in causal dynamics and generalizes effectively to previously unseen contexts.
@@ 1 Introduction
To identify the underlying world rules, traditional causal models rely predominantly on observational data (pmlr-v216-lippe23a), implicitly assuming a fixed causal structure. In open‐ended environments, although the ground‐truth causal laws remain unchanged, the observational data is collected through local, context‐limited exploration. Under conditions of partial observability, this invariance becomes obscured within a limited observational window, giving rise to the illusion of shifting causal relationships—a rule valid in one context may break down in another. For instance, the "push → open" relation holds for an unlocked door, but this relation vanishes when the door is locked. In this case, the same action produces no effect, leaving the true causal mechanism ambiguous. Traditional causal modeling in such scenarios faces two main challenges: (1) methods that learn a single uniform (time series-based) causal graph neglect these context‐dependent changes (NEURIPS2023_65496a49), and (2) lack of an active curiosity-driven exploration strategy, which makes it struggle to adapt to discover the global causal rules in rich, unseen variations of potential open-ended worlds (huang2020causal; hwang2024fine).
@@ 2 Preliminaries
Causal Graph.
A causal subgraph is formalized as a directed acyclic graph (DAG) $\mathcal{G}=([p],E)$, where the vertex set $p]=\{1,2,\dots,p\}$ indexes a collection of random variables $X=\{X_{i}\}_{i=1}^{p}$ with joint density $f(X)$. The edge set $E\subseteq[p]\times[p]$ encodes direct causal relationships between variables. For each node $i\in[p]$, we write $\mathrm{Pa}_{\mathcal{G}}(i)$ for its set of parents. An edge $i\to j\in E$ indicates that $X_{i}$ exerts a direct causal influence on $X_{j}$.
Intervention.
@@
During world‐model learning, we perform a sequence of environment interventions to reveal causal structure. To better formalize the intervention influence, in this paper, we introduce the notion of an intervention target. At intervention step $k$, we select a single variable $X_{i}$ as the target, which we denote by the intervention target $I_{k}=\{i\}$, where $i\in[p]$. Applying an intervention $\mathrm{do}(X_{I_{k}}=x^{\prime})$ removes all incoming causal links into $X_{i}$. Concretely, if $\mathcal{G}$ is the original DAG, the intervention graph $\mathcal{G}^{(k)}$ is obtained by removing every edge of the form $j\to i,\quad\forall\,j\in\mathrm{Pa}_{\mathcal{G}}(i),$ from $\mathcal{G}$. This modified graph accurately reflects the altered generative process under the do‐operation on $X_{i}$.
@@
Let $D=([p],E)$ be a DAG with vertex set $[p]$ and edge set $E$, and $I\subset[p]$ an intervention target. The intervention graph of $D$ is the DAG $D^{(I)}=([p],E^{(I)})$, where $E^{(I)}:=$ $\{a\to b\mid a\to b\in E,b\notin I\}.$
@@ 3 Meta-Causal Graph: Definition and Identifiability
In this section, we present our Meta‐Causal Graph for a world formulation. We begin by defining the Meta‐Causal Graph, which is a compact causal graph comprising multiple context‐specific causal subgraphs, varying from latent meta states in Section 3.1. Then, we show the identifiability of meta states in Section 3.2 and their indicated causal subgraphs in Section 3.3 by interventions.
@@ 3.1 Meta-Causal Graph
The Meta‐Causal Graph $\mathcal{MG}$ consists of a collection of causal subgraphs $\mathcal{MG}\;=\;\{\mathcal{G}_{u}\}_{u\in U},$ where each $\mathcal{G}_{u}$ corresponds to a distinct meta state $u\in U$. The causality skeleton matrix $M_{u}$ is a binary matrix where $M_{u}[i,j]=1$ indicates that variable $i$ is a parent of variable $j$ (i.e., $i\in Pa(j)$) in the causal graph $\mathcal{G}_{u}$. A state‐to‐meta state mapping $C:\mathcal{X}\;\to\;U$ assigns each system observation $\mathbf{x}\in\mathcal{X}$ to its active meta state $u=C(\mathbf{x})$. We call $\mathcal{MG}=(\{\mathcal{G}_{u}\},C)$ a Meta‐Causal Graph if the following conditions are satisfied:
@@
    Causality skeleton matrices of causal subgraphs are sufficiently different. For all $u\neq u^{\prime}$, $M_{u}\neq M_{u^{\prime}}$, i.e., there exists an index $(i,j)$ such that $M_{u}[i,j]\neq M_{u^{\prime}}[i,j]$.
@@ 3.2 Identifiability of Meta States
Intuitively, observational equivalence refers to the condition where two mappings from states to meta states produce the same partitioning of observations, possibly up to relabeling.
@@ 3.3.1 Causal Subgraph Identification
We employ multiple interventions to determine the causal structure of environment, represented as intervention target indicator set $\mathcal{I}=\{I_{k}\}_{k=1}^{K}$, where $I_{k}\subset\{[p]\}$ comprises the indices of intervened variables selected from state variable set $X=\{X_{i}\}_{i\in[p]}$ at intervention step $k$. These intervention sets are collected from historical data or agent interactions. We establish sufficient conditions ensuring causal graph edge identifiability through appropriate interventional selection.
@@
4 Curious Causality-Seeking Agents for Open-Ended World Modeling
@@
We propose a novel framework to learn the Meta-Causal Graph through a Causality-Seeking Agent. Our approach comprises three core components: (1) curiosity-driven interventional exploration in an open-ended world, (2) Meta-Causal Graph discovery from agent experience, and (3) continual world model learning and updating. Together, these components enable the agent to actively discover and refine causal structures in open-ended environments.
4.1 Curiosity-Driven Interventional Exploration in Open-Ended Worlds
Accurate identification of the Meta-Causal Graph cannot rely solely on passive experience. First, causal structures inferred from purely observational trajectories often suffer from edge misorientation. Second, an agent’s experience typically spans only a limited subset of world states, leaving parts of the causal graph unobserved. To reveal the full Meta-Causal Graph in an open-ended environment, the agent must complement passive observation with targeted, curiosity-driven interventions that actively probe uncertain causal relations.
Curiosity Reward.
Our framework treats curiosity as a general intrinsic signal that can be instantiated in multiple forms. Specifically, we experimented with the following formulations:
(1) Edge-Entropy. $I^{\text{edge}}_{t}=\sum_{i,j\in[p]}H\big(\hat{M}_{C(X_{t})}[i,j]\big),$ where $\hat{M}_{C(X_{t})}[i,j]\!\in[0,1]$ denotes the posterior probability of an edge from variable $i$ to $j$ in the current meta state $C(X_{t})$, and $H(\cdot)$ is the Shannon entropy. This term prioritizes interventions on the most uncertain parts of the causal graph.
(2) Prediction Uncertainty. $\mathcal{I}^{\text{unc}}_{t}=H\big[p_{\theta}(x_{t+1}\mid x_{t},a_{t})\big],$ where $p_{\theta}(x_{t+1}\mid x_{t},a_{t})$ is the world model’s predictive distribution over the next state given current state–action pair $(x_{t},a_{t})$. Higher entropy indicates epistemic uncertainty about future transitions.
(3) Feature Discrepancy. $\mathcal{I}^{\text{feat}}_{t}=\big\|E(x_{t+1})-E_{\theta}(\hat{x}_{t+1})\big\|_{2}^{2},$ where $\phi(\cdot)$ is the learned feature encoder and $E_{\theta}(\hat{x}_{t+1})$ is the predicted feature of the next state $\hat{x}_{t+1}$. This term measures reconstruction error in the feature space, highlighting dynamics that the model fails to capture.
(4) Predictive-Distribution Discrepancy. $\mathcal{I}^{\text{nll}}_{t}=-\log p_{\theta}(x_{t+1}\mid x_{t},a_{t}),$ which is the surprisal or negative log-likelihood of the observed transition.
Any of these intrinsic terms can be used as the curiosity reward $R_{t}$ to select interventions, and in our main results we adopt the edge-entropy variant as the default while other forms yield comparable behaviors (see Table 9).
Intervention Verification.
After executing curiosity-guided interventions, we validate and refine the learned causal structures by directly estimating causal effects. For each variable $X_{i}$, we perform interventions $\mathrm{do}(X_{i}{=}x^{\prime}_{i})$ and observe the resulting changes in other variables. We estimate the causal effect of $X_{i}$ on $X_{j}$ as
$\Delta_{ij}=\log P(X_{j}^{t+1}\mid X^{t},\mathrm{do}(X_{i}{=}x^{\prime}_{i}))-\log P(X_{j}^{t+1}\mid X^{t}).$
We integrate these effects into a mask-refinement loss
$\mathcal{L}_{\text{mask}}=-\lambda_{1}\!\!\sum_{\{(i,j):|\Delta_{ij}|>\tau\}}\!\!\log\hat{M}_{ij}+\lambda_{2}\!\!\sum_{\{(i,j):|\Delta_{ij}|<\tau\}}\!\!\log\hat{M}_{ij},$
where $\tau\geq 0$ controls sensitivity.
Interventional Reachability.
In realistic environments, not all variables are directly intervenable. To formalize which causal relations are identifiable under such constraints, we introduce interventional reachability. We define an intervention operator $F\in\{0,1\}^{N\times N}$ where $F[k,i]=1$ if and only if state $k$ can be obtained from state $i$ by a single allowed intervention, and a transition operator $T\in\{0,1\}^{N\times N}$ for the environment’s natural dynamics after intervention. The composite $TF$ models one intervention–transition cycle, and nonzero entries of $(TF)^{k}z_{0}$ enumerate states reachable within $k$ cycles. A state is reachable if and only if some finite $k$ satisfies $(TF)^{k}z_{0}][i]>0$. These constraints define the feasible set of causal interventions. Details are given in Appendix E.
4.2 Meta-Causal Graph Discovery from Agent Experience
Given trajectories collected through curiosity-driven interventions, the agent infers a structured representation of causal dependencies that vary across latent meta states. The goal is to jointly learn (i) the discrete latent variable $C(X_{t})$ indicating the current meta state, and (ii) the corresponding causal skeleton matrix $M_{C(X_{t})}$ that governs the local dynamics.
Representation Learning.
To model context-dependent causal dependencies from trajectories generated by curiosity-driven interventions, we discretize the latent dynamics of the environment using a vector-quantized variational auto-encoder (VQ-VAE) (hwang2024fine). The encoder maps concatenated state–action pairs $(x_{t},a_{t})$ into a latent vector $z_{e}$, which is quantized to the nearest codebook entry $z_{u}$ associated with a meta state $u$. Each codebook entry defines a distribution over causal structures, from which the causal skeleton matrix $M_{C(X_{t})}$ is sampled. Further details are provided in Appendix E.
Effect of Lossy Representation.
The latent representation learned through vector quantization is inherently lossy due to imperfect mapping. Such lossy representations may distort the underlying causal factors and thus influence the identification of both meta states and their associated causal subgraphs. We provide an error bound of lossy representation in the following.
Proposition 2 (Effect of Representation Accuracy on Misclassification).
Given a set of ground-truth meta states $U=\{u\}$ and corresponding codebook entry index $\hat{U}=\{\hat{u}_{k}\}$, let $p_{k}=\sum_{u\in U}\mu_{u}p_{k}^{u}$ denote the probability that a sample $x$ is mapped to codebook entry $z_{\hat{u}_{k}}$. Then, under the lossy representation induced by encoder, the probability of misclassification is
$P(\text{misclassification})=1-\sum_{k}\sum_{u\in U}\mu_{u}p^{u}_{k}(1-p_{k}+p^{u}_{k}\mu_{u})^{n}.$
When $n(p_{k}+p^{u}_{k}\mu_{u})\ll 1$, this can be approximated as
$P(\text{misclassification})\approx\sum_{k}\left[p_{k}^{2}-\sum_{u}(\mu_{u}p^{u}_{k})^{2}\right].$
Adaptive Codebook Fusion.
To prevent redundant meta states resulted from overparamterization and maintain a minimal representation, we adopt an Adaptive Codebook Fusion mechanism. During training, codebook entries encoding similar latent embeddings and producing comparable decoded causal skeleton matrices are automatically merged. This fusion step is integrated into the quantization update, requires no extra supervision, and reallocates capacity for novel meta states.
@@ 4.3 Transition Probability Learning
[…]
@@
1: Initialize encoder parameters $\phi$, decoder parameters $\theta$, and embeddings $\{z_{u}\}_{u\in U}$
@@
3:  Compute embedding assignment: $u\leftarrow\operatorname*{arg\,min}_{u\in U}\|E_{\phi}(X)-z_{u}\|_{2}^{2}$
@@
6:  Select intervention $\text{do}(X_{i}=x^{\prime}_{i})$ maximizing reward $R_{t}$ in Section 4.1